OpenStudy (anonymous):

int (arctan(e^x))/(1+e^2x)dx

hartnn (hartnn):

did you try the substitution u = e^x ? can you also verify your question ? because the funtion you posted is not integrable in terms of standard math functions.

hartnn (hartnn):

i mean, try e^x = tan y

OpenStudy (anonymous):

i did try letting u=e^x. it did not get me anywhere . The limits on te question are really from 0 too 2 and the question is asking for the numerical value \[\int\limits_{0}^{?}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{2}(\arctan(e^x))/(1+e^2x)dx\]

hartnn (hartnn):

\(\large \int\limits_{0}^{2}\dfrac {(\arctan(e^x))}{(1+e^{2x})}dx\) with e^x = tan y e^x dx = sec^2 y dy dx = sec^2 y/ tan y dy

OpenStudy (anonymous):

how is e^x dx = sec^2 y dy

hartnn (hartnn):

took derivative on both sides derivative of e^x is e^x dx derivative of tan y is sec^2 y dy

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

\[\frac{ \tan y }{-\ln \cos y }\] Am i on the right track ?

hartnn (hartnn):

don't you just get y/tan y ?